OpenStudy (anonymous):

(X^3/4)^1/5 Please help simplify this

OpenStudy (jdoe0001):

\(\bf \Large \left(x^{\frac{3}{4}}\right)^{\frac{1}{5}}\implies \huge{x}^{\frac{3}{4}\times\frac{1}{5}}\)

OpenStudy (anonymous):

ok so do i cross simplify now ?

OpenStudy (jdoe0001):

well, is a multiplication

OpenStudy (jdoe0001):

\(\bf \large\left(x^{\frac{3}{4}}\right)^{\frac{1}{5}}\implies x^{\frac{3}{4}\cdot\frac{1}{5}}\)

OpenStudy (anonymous):

oh ok i see Thankyou !!!!!!!!!!!

OpenStudy (jdoe0001):

yw